OpenStudy (anonymous):

Harold left town driving east at 55 mph. Carol left 2 hours later driving the same route at 65 mph. How many hours after Harold left will it take carol to catch up to Harold? A) 10 B)11 C)12 D)13

OpenStudy (anonymous):

55t = 65(t-2) 55t = 65t-130 130 = 10t t = 13

OpenStudy (anonymous):

:D